### v1

@@ 2.1 Preliminaries
In this study, we focus on optimizing interventions for individual process instances in isolation, an assumption that simplifies the problem by disregarding interdependencies between cases. In real-world scenarios, such interdependencies—often due to shared resource constraints—can significantly impact outcomes. Recent research has begun to address these complexities [25, 31], but they remain outside the scope of this work.
@@ 2.2 Prescriptive Process Monitoring Approaches
The first major approach adopted in PresPM is Causal Inference (CI), which focuses on estimating the effect of an intervention action on a target variable for each individual case using offline data, whether synthetic or real [13]. A core challenge in CI is that only the outcome of the chosen intervention action is recorded for each case, while the counterfactual outcome remains unknown. One way to estimate this counterfactual outcome is to analyze similar cases in the data that received a different intervention action. However, real-world data-gathering policies often introduce confounding, where other variables influence both treatment/intervention and outcome. This can lead to treated and untreated cases differing strongly in their distributions, unlike the balanced and confounding-free setup of RCTs. Traditional CI methods attempt to correct for such distortions [18]. An example of CI in PresPM is the method described in [9], which incorporates CI by augmenting logs with estimated counterfactuals to identify optimal causal effect estimators for interventions. RealCause, a generative AI framework originally used as a CI evaluation tool, facilitates this process. Other studies include [29], [30], [3], and [31], although CI is not necessarily the main focus in the latter two.
@@ 2.3 Limitations in Prescriptive Process Monitoring
Current approaches in PresPM mainly rely on historical offline datasets, with BPIC12 and BPIC17 [10, 11] being by far the most commonly used in works such as [9, 3, 31, 4]. Although insightful, historical datasets lack counterfactual outcomes. While we can estimate the impact of interventions, their true effects remain unknown, resulting in only approximate evaluations. For example, in [3], RealCause not only generates alternative outcomes for the RL environment, but also defines the evaluation metrics, assuming that these generated outcomes are the true outcomes. Furthermore, comparison studies of PresPM methods are scarce. To the best of our knowledge, [34] is the only work to compare two implementations comprehensively by using a synthetic setup for precise evaluation. This scarcity is partly due to the lack of accurate evaluation mechanisms and unclear intervention dimensions, making comparisons difficult. These dimensions, i.e., action width and depth, are crucial to know, as the effectiveness of PresPM methods depends on them. For example, [34] shows that their CI implementation, though safer, performs worse than RL in scenarios with action depths larger than 1 (timing choice). It is also worth noting that most research focuses on interventions with action width 2 and action depth larger than 1 [9, 30, 31]. For example, in BPIC12 and BPIC17, the act of sending multiple offers is considered an intervention and follows these dimensions. Optimization of intervention sequences is underexplored. Only the methods in [33] and [17] address this by deploying multiple predictive models. However, their approaches may accumulate errors, leaving room for improvement. Lastly, commonly used logs often lack empirically validated causal relationships. Researchers typically select interventions without external domain expert confirmation, especially regarding their heterogeneous effects across different cases. For instance, while the intervention in BPIC12 and BPIC17 is correlated with higher loan acceptance rates, its causal impact has not been extensively studied or externally validated.
Simulators offer a solution to these limitations. In a simulator, the process-generative function is fully known, allowing for the calculation of outcomes for all possible interventions, thus supporting accurate evaluation by generating counterfactuals. This setup also enables controlled experimentation, e.g., parameters such as confounding level can be adjusted systematically. A simulator provides the flexibility to clearly define a range of interventions with varying complexity, and explicitly model their causal effects on outcomes. Combining these interventions also supports research for intervention sequences.
@@ 2.4 Current Simulations in Process Mining
Given the benefits of a custom simulator designed for PresPM, it is important to understand why a new simulator may be preferred over existing simulators. While CI and RL studies outside of Process Mining often use simulations [2, 32], these do not generate business process data. Simulators have also gained traction within Process Mining. For example, Data-Driven Process Simulation (DDPS) approaches include [7, 14, 26]. Others train and employ DL models to generate simulated event logs, such as in [6]. Examples of approaches that combine both, or hybrid simulations, are [5, 20]. Still, existing Process Mining simulators are not designed with a PresPM focus. Firstly, these simulations are based on historical datasets, and thus still suffer from the same counterfactual limitations. Secondly, most simulators are not designed to study interventions, as they lack a well-defined, outcome-influencing intervention mechanism. Additionally, they do not support multiple intervention types or allow controlled experiments relevant to PresPM research, such as manipulating confounding levels. While [34] does use a simulation to evaluate PresPM methods, the simulated process is overly simplistic, lacking loops, parallel flows, and diverse intervention types. This lack of suitable simulation setups in PresPM, compared to CI, is likely due to the challenge of modeling (non-trivial) time dependencies between variables, control-flow, interventions, and outcomes. Nevertheless, a purpose-built simulator tailored to PresPM research could significantly help in addressing these current limitations. By supporting complex interventions and realistic process features, it can provide a foundation for systematic evaluation, experimentation, and the exploration of intervention sequences in PresPM.
@@ 3 Simulator Development
In this section, we introduce the development of SimBank. We begin by outlining the high-level requirements derived from our review of the literature, followed by a description of the process and its interventions. We then delve into the implementation details, highlighting SimBank’s capabilities. Finally, we discuss the generality of SimBank.
@@ 3.2 Description of the Process and Interventions
The design of the simulator’s process model centers on a fictional loan application process in a bank. The model is illustrated in Figure 1, which includes activity costs, durations, the bank’s loan processing policy, and external environmental factors. The process begins with an application submission, followed by the bank selecting a procedure, potentially contacting HQ and the customer, calculating an interest rate, or canceling the application. If an interest rate is offered, the client’s response follows. Table 2 lists the data attributes guiding the bank policy and the client’s (actual but unobservable) quality, an external factor beyond the bank’s control. The target variable to optimize is the loan profit, considering variables like the amount requested, interest rate, time elapsed, costs, (actual) client quality, and cancellation reasons. The client’s decision also affects profit, but is in itself influenced by elapsed time, interest rate, and amount. For simplicity, potential client default is excluded. We have integrated three main interventions into the loan application process: Choose procedure, Set interest rate, and Time contact HQ. The intervention dimensions and details are outlined in Table 3.
@@ 3.3 Implementation and Capabilities
The simulator is built in Python, primarily using the simpy library [22]. While the initial code is based on the PNSIM framework [26], it is extensively modified to suit the requirements of our research, incorporating the specific features discussed below. We use random number generators to generate variables. The code and comprehensive guidance on using SimBank can be found at https://anonymous.4open.science/r/SimBank-E3FA.
@@
3.4 Notes on the Generality of SimBank.
A more generalized simulation approach, applicable to any event log, would be preferable to SimBank. However, as noted earlier, simulations only based on historical data lack verified causal dependencies and result in approximate evaluations. Accurate evaluations require specifying these dependencies, which is complex, time-consuming, and process-specific. This makes a log-agnostic simulation approach for PresPM largely ineffective as it would still require users to manually define the causal relations over time for each log. For instance, fully accurate synthetic evaluations in BPIC17 would require defining which clients receive multiple offers (the intervention) and when, as well as modeling how the application evolves when an intervention has taken place, or not (yet). Each activity essentially requires a function that takes case variables (e.g., loan amount), activity history, event variables (e.g., bank resources), and the intervention action, and outputs the next activity and any changes to event variables. At the case’s end, a function must determine the outcome based on these factors as well. Most importantly, all these time-sensitive causal dependencies must be case-specific and non-trivial to differentiate between PresPM methods. This would still result in just one defined intervention, rather than multiple, as seen in SimBank.
Consequently, although SimBank is based on a single process, its generality is ensured in other ways. First, as mentioned, SimBank supports multiple intervention types across various dimensions, each with explicitly defined (heterogeneous) causal effects, providing more versatility and clarity for research compared to existing logs (in addition to the accuracy of the evaluation). Second, as stated in the previous subsection, users can modify the bank policy or add new activities and interventions, expanding or fully adapting the process through the simulator’s functions. Lastly, we argue that SimBank has generality as it can serve as an independent evaluation setup: its development is separate from that of any PresPM method, ensuring that when used in a study, there is no risk of synthetic experiments favoring the proposed method—which can occur when solution and evaluation are developed within the same work.
@@ 4.1 Experimental Setup
We experiment with 4 PresPM methods—2 based on CI and two based on RL—to assess their effectiveness in optimizing interventions. We also use the RealCause evaluation method to compare its results with the accurate SimBank evaluation. First, we explain the RealCause implementation, followed by the 4 PresPM methods.
  • Evaluation method: RealCause. RealCause was developed to benchmark CI approaches by generating counterfactuals that reflect a real dataset, helping to assess whether a CI method recommends the optimal actions. RealCause has been used in PresPM research, such as in [3, 9]. Given the intervention action recommended by the PresPM approach for a test case, RealCause generates the corresponding outcome. The method models two key distributions: the selection distribution, which determines which intervention action is assigned to a case and is modeled here by a Bernoulli distribution, and the outcome distribution, modeled by a Sigmoid Flow distribution with discrete value concentrations [12]. The distributions are approximated using a TARNet-based structure, which includes a neural network to learn treatment-agnostic representations and two sub-networks tailored to specific treatment groups [28]. This design could be effective as it utilizes all data for the initial representation network while emphasizing treatment through the separate networks. The models were tested using the statistical tests of the original RealCause paper [24].
@@
All methods are trained on 100,000 cases, with 1,000 for validation and 10,000 for testing. RealCause and the K-means model are trained on an extra unrelated dataset of 10,000 cases. The evaluation metric, denoted as Gain, is the relative change in total profit compared to the bank’s baseline policy. Each experiment is repeated 5 times.
@@ 4.2 Results
Table 4 presents the results for the Choose procedure and Set interest rate interventions. In both cases, the standard S-learner and Deep Q-learning outperform the random policy and bank policy. For the Choose procedure intervention, the standard S-learner demonstrates the best performance and greater stability, with a lower standard deviation. For the Set interest rate intervention, the standard S-learner also slightly outperforms Deep Q-learning and remains the more stable method. While Q-learning theoretically offers optimal solutions [32], its practical implementation underperforms. Potential improvements include exploring alternative state representations, architectures, or efficient sampling techniques [19].
@@
Table 6 presents results for the Time contact HQ intervention with $\delta=0.999$, creating a heavily confounded training dataset. All methods outperform the random and bank policies, though the standard S-learner only marginally exceeds the bank policy. True evaluation shows the RealCause-based S-learner performs best, while the standard S-learner is more stable. However, performance suffers across the board, as the dataset is confounded and lacks diversity, omitting some optimal actions. The RealCause-based S-learner performs better, likely due to the TARNet-based structure in RealCause which emphasizes treatment selection. This suggests that using TARNet directly could further improve performance. Importantly, the RealCause evaluation model, trained on a confounded dataset, overestimates the RealCause-based S-learner and random policy, while underestimating the standard S-learner. The alignment between the RealCause-based S-learner and the RealCause evaluation model inflates its performance. These findings highlight the risks of relying on RealCause for both data augmentation and evaluation, as seen in [3], and emphasize the need for accurate evaluations, as evaluation models like RealCause can lead to inaccuracies due to confounding.
@@
Figure 2 contains the results of the experiments on the standard S-learner where we vary $\delta$ at 6 different values. The minimum value of $0.95$ is chosen since performance stays more or less consistent between 00 and $0.95$ for a dataset of 100,000 cases. The error bars indicate a confidence interval of the performance at each level. Except for the Set interest rate intervention, a general trend is that both performance and stability decrease for higher values of $\delta$. Especially the bank’s original policy for the Time contact HQ intervention seems suboptimal, given the high sensitivity of the standard S-learner to $\delta$. For the sequence of Choose procedure and Set interest rate, sequential optimization proves beneficial, as our setup outperforms optimizing only one of the interventions. However, since our implementation consists of one model per intervention, each model optimizes its intervention independently, assuming the bank policy for the other. This likely limits overall performance. More research is needed to improve the optimization of intervention sequences, which is especially relevant in business processes.
@@ 5.0.1 Acknowledgments
This work was supported by the Research Foundation - Flanders (FWO research projects 11A6J25N and G039923N).

### v2

@@ 2.1 Preliminaries
In this study, we focus on optimizing interventions for individual process instances in isolation, an assumption that simplifies the problem by disregarding interdependencies between cases. In real-world scenarios, such interdependencies—often due to shared resource constraints—can significantly impact outcomes. Recent research has begun to address these complexities [25, 31], but they remain outside the scope of this work—a simplification also made in the majority of existing studies.
@@ 2.2 Prescriptive Process Monitoring Approaches
The first major approach adopted in PresPM is Causal Inference (CI), which focuses on estimating the effect of an intervention action on a target variable for each individual case using offline data, whether synthetic or real [13]. A core challenge in CI is that only the outcome of the chosen intervention action is recorded for each case, while the counterfactual outcome remains unknown. One way to estimate this counterfactual outcome is to analyze similar cases in the data that received a different intervention action. However, real-world data-gathering policies often introduce confounding, where other variables influence both treatment/intervention and outcome. This can lead to treated and untreated cases differing strongly in their distributions, unlike the balanced and confounding-free setup of RCTs. Traditional CI methods attempt to correct for such distortions [18]. An example of CI in PresPM is the method described in [9], which incorporates CI by augmenting logs with estimated counterfactuals to identify optimal causal effect estimators. RealCause, a generative AI framework originally used as a CI evaluation tool, facilitates this process. Other studies include [29], [30], [3], and [31], although CI is not necessarily the main focus in the latter two.
@@ 2.3 Limitations in Prescriptive Process Monitoring
Current approaches in PresPM mainly rely on historical offline datasets, with BPIC12 and BPIC17 [10, 11] being by far the most commonly used in works such as [9, 3, 31, 4]. Although insightful, historical datasets lack counterfactual outcomes. While we can estimate the impact of interventions, their true effects remain unknown, resulting in only approximate evaluations. RealCause is the most widely used evaluation method to obtain these approximate evaluations [9, 3, 31, 30, 32]. For example, in [3], RealCause not only generates alternative outcomes to train an RL agent, but also defines the evaluation metrics, assuming that these generated outcomes are the true outcomes. Furthermore, comparison studies of PresPM methods are scarce. To the best of our knowledge, [35] is the only work to compare two implementations comprehensively by using a synthetic setup for precise evaluation. This scarcity is partly due to the lack of accurate evaluation mechanisms and unclear intervention dimensions, making comparisons difficult. These dimensions, i.e., action width and depth, are crucial to know, as the effectiveness of PresPM methods depends on them. For example, [35] shows that their CI implementation, though safer, performs worse than RL in scenarios with action depths larger than 1 (timing choice). It is also worth noting that most research focuses on interventions with action width 2 and action depth larger than 1 [9, 30, 31]. For example, in BPIC12 and BPIC17, the act of sending multiple offers is considered an intervention and follows these dimensions. Optimization of intervention sequences is underexplored. Only the methods in [34] and [17] address this by deploying multiple predictive models. However, their approaches may accumulate errors, leaving room for improvement. Lastly, commonly used logs often lack empirically validated causal relationships. Researchers typically select interventions without external domain expert confirmation, especially regarding their heterogeneous effects across different cases. For instance, while the intervention in BPIC12 and BPIC17 is correlated with higher loan acceptance rates, its causal impact has not been extensively studied or externally validated.
Manually defined simulators offer a solution to these limitations. In such a simulator, the process-generative function is fully known by design, allowing for the calculation of outcomes for all possible interventions, thus supporting accurate evaluation by generating counterfactuals. This setup also enables controlled experimentation, e.g., parameters such as confounding level can be adjusted systematically. Fully specifying the simulator from the ground up provides the flexibility to clearly define a range of interventions with varying complexity within the same process, and explicitly model their causal effects on outcomes. Combining these interventions also supports research for intervention sequences.
@@ 2.4 Current Simulations in Process Mining
Given the benefits of a manually defined simulator for PresPM, it is important to understand why such a new simulator may be preferred over existing simulators. While CI and RL studies outside of Process Mining often use simulations [2, 33], these do not generate business process data. Simulators have also gained traction within Process Mining. For example, Data-Driven Process Simulation (DDPS) approaches include [7, 14, 26]. Others train and employ DL models to generate simulated event logs, such as in [6]. Examples of approaches that combine both, or hybrid simulations, are [5, 20]. Extending these simulations to PresPM research would be useful, as they can be applied to any event log and could enable realistic experiments. However, they have a major limitation in this context: these simulations are based on historical datasets, and thus still suffer from the same counterfactual limitations. They are typically developed to reflect the policy that generated the original event log (e.g., bank policy), making them unlikely to accurately generate outcomes when a PresPM method recommends actions that differ from past practice. Additionally, they do not support multiple intervention types or allow controlled experiments relevant to PresPM research, such as manipulating confounding levels. Even if certain components are manually defined to extend these simulations, the evaluation accuracy remains uncertain due to the limitations in the parts that are not manually defined but approximated.
In contrast, manually defined simulators provide the opposite trade-off: they offer accuracy in evaluation and enable fully customizable experiments, including support for multiple interventions and confounding scenarios. However, they are not automatically generalizable to any event log and may lack realism. Currently, only one such simulator exists for PresPM [35], but it is overly simplistic, lacking loops, parallelism, and a variety of interventions. This reveals a clear gap and opportunity for developing a more sophisticated, realistic specified simulator tailored to PresPM needs that provides customizable and fully accurate experiments.
@@ 3 Simulator Development
In this section, we introduce the development of SimBank. We begin by outlining the high-level requirements derived from our review of the literature, followed by a description of the process and its interventions. We then delve into the implementation details, highlighting SimBank’s capabilities.
@@ 3.2 Description of the Process and Interventions
The design of the simulator’s process model centers on a fictional loan application process in a bank. The model is illustrated in Figure 1, which includes activity costs, durations, the bank’s loan processing policy, and external environmental factors. The process begins with an application submission, followed by the bank selecting a procedure, potentially contacting HQ and the customer, calculating an interest rate, or canceling the application. If an interest rate is offered, the client’s response follows. Table 2 lists the data attributes guiding the bank policy and the client’s (actual but unobservable) quality, an external factor beyond the bank’s control. The target variable to optimize is the loan profit, considering variables like the amount requested, interest rate, elapsed time, costs, (actual) client quality, and cancellation reasons. The client’s decision also affects profit, but is in itself influenced by elapsed time, interest rate, and amount. For simplicity, potential client default is excluded. We have integrated three main interventions into the loan application process: Choose procedure, Set interest rate, and Time contact HQ. The intervention dimensions and details are outlined in Table 3.
@@ 3.3 Implementation and Capabilities
The simulator is built in Python, primarily using the simpy library [22]. While the initial code is based on the PNSIM framework [26], it is extensively modified to suit the requirements of our research, incorporating the specific features discussed below. We use random number generators to generate variables. The code and comprehensive guidance on using SimBank can be found at https://github.com/JakobDeMoorKULstudent/SimBank.
@@ 4.1 Experimental Setup
We experiment with 4 PresPM methods—2 based on CI and two based on RL—to assess their effectiveness in optimizing interventions. We also use the widely adopted RealCause evaluation method to compare its approximated evaluation results with the accurate SimBank evaluation. First, we explain the RealCause implementation, followed by the 4 PresPM methods.
  • Evaluation method: RealCause. RealCause was developed to benchmark CI approaches by generating counterfactuals that reflect a real dataset, helping to assess whether a CI method recommends the optimal actions. RealCause has been used in PresPM research, such as in [9, 3, 31, 30, 32]. Given the intervention action recommended by the PresPM approach for a test case, RealCause generates the corresponding outcome. The method models two key distributions. The first is the selection distribution, which determines which intervention action is assigned to a case and is modeled here by a Bernoulli distribution for interventions Choose procedure and Time contact HQ, and a categorical distribution for Set interest rate. The second is the outcome distribution, modeled by a Sigmoid Flow distribution with discrete value concentrations [12]. The distributions are approximated using a TARNet-based structure, which includes a neural network to learn treatment-agnostic representations and two sub-networks tailored to specific treatment groups [28]. This design could be effective as it utilizes all data for the initial representation network while emphasizing treatment through the separate networks. The models were tested using the statistical tests of the original RealCause paper [24].
@@
All methods are trained on 100,000 cases, with 1,000 for validation and 10,000 for testing. RealCause and the K-means model are trained on an extra unrelated dataset of 10,000 cases. The evaluation metric, denoted as Gain, is the relative change in total profit compared to the bank’s baseline policy. A policy therefore outperforms the bank policy if its $\textit{Gain}>0$. Each experiment is repeated 5 times.
@@ 4.2 Results
Table 4 presents the results for the Choose procedure and Set interest rate interventions. Note that the performance of the bank policy is not in the table, as its Gain is naturally 0. In both cases, the standard S-learner and Deep Q-learning outperform the random policy and bank policy. For the Choose procedure intervention, the standard S-learner demonstrates the best performance and greater stability, with a lower standard deviation. For the Set interest rate intervention, the standard S-learner also slightly outperforms Deep Q-learning and remains the more stable method. While Q-learning theoretically offers optimal solutions [33], its practical implementation underperforms. Potential improvements include exploring alternative state representations, architectures, or efficient sampling techniques [19].
@@
Table 6 presents results for the Time contact HQ intervention with $\delta=0.999$, creating a heavily confounded training dataset. Note that the performance of the random policy remains unchanged from the unconfounded setting, since it does not rely on the training data. All methods outperform the random and bank policies, though the standard S-learner only marginally exceeds the bank policy. True evaluation shows the RealCause-based S-learner performs best, while the standard S-learner is more stable. However, performance suffers across the board, as the dataset is confounded and lacks diversity. The RealCause-based S-learner performs better, likely due to the TARNet-based structure in RealCause which emphasizes treatment selection. Moreover, the RealCause evaluation model leads to inaccurate evaluations when trained on a confounded dataset. It overestimates the RealCause-based S-learner, due to the alignment between this learner and the RealCause evaluation model. This underscores the risk of relying on RealCause for both data augmentation during training and for defining evaluation metrics, as seen similarly in [3], albeit with an RL agent instead. More critically, RealCause ranks the random policy higher than the standard S-learner, despite the true evaluation indicating otherwise. These findings emphasize the importance of accurate and unbiased evaluation methods, such as those provided by SimBank. Since most real-world event logs are collected under existing, non-randomized policies (unlike RCTs), they are likely to be confounded. In such cases, relying solely on RealCause can lead to misleading conclusions.
@@
Figure 2 contains the results of the experiments on the standard S-learner where we vary $\delta$ at 6 different values. The minimum value of $0.95$ is chosen since performance stays more or less consistent between 00 and $0.95$ for a dataset of 100,000 cases. The error bars indicate a confidence interval of the performance at each level. Except for the Set interest rate intervention, a general trend is that both performance and stability decrease for higher values of $\delta$. Especially the bank’s original policy for the Time contact HQ intervention seems undermining, given the high sensitivity of the standard S-learner to $\delta$. For the sequence of Choose procedure and Set interest rate, sequential optimization proves beneficial, as our setup outperforms optimizing only one of the interventions. However, since our implementation consists of one model per intervention, each model optimizes its intervention independently, assuming the bank policy for the other. This likely limits overall performance. More research is needed to improve the optimization of intervention sequences, which is especially relevant in business processes.
@@ 5.0.1 Acknowledgments
This work was supported the Research Foundation Flanders (FWO) under grant number G039923N, and Internal Funds KU Leuven under grant number C14/23/031.